# MIRROR: Multi-Modal Pathological Self-Supervised Representation Learning via Modality Alignment and Retention

### v2

@@ IV-D Ablation Study
Alignment Retention Clustering Subtyping Survival
0.953 $\pm$ 0.024 0.612 $\pm$ 0.075
0.954 $\pm$ 0.028 0.610 $\pm$ 0.085
0.950 $\pm$ 0.016 0.654 $\pm$ 0.083
0.955 $\pm$ 0.023 0.665 $\pm$ 0.082
We conducted an ablation study to evaluate the effectiveness of the key components of MIRROR, including the modality alignment module, the modality retention module, and the style clustering module. The study was performed on the TCGA-BRCA dataset using linear probing with Phikon as the backbone, comparing MIRROR’s performance both with and without these components.
As shown in Table IV, the results highlight the contributions of each module. The modality retention module, while leading to a slight decline in survival analysis performance, enhances subtype classification, demonstrating its ability to preserve modality-specific information. Additionally, the introduction of the style clustering module encourages the model to focus on disease-relevant patterns while mitigating redundancy, enabling the extraction of more informative features. This refinement is reflected in the improved feature retention performance observed during training. With all modules incorporated, MIRROR achieves the best overall performance, confirming the synergistic benefits of the proposed components.
Here, we demonstrate the explainability of MIRROR with attention weights visualization from the slide encoder and UMAP [62] projections of features extracted by RNA encoder.
As illustrated in Figure 4, we selected four representative samples from each cohort used, respectively. The attention weights from the slide encoder are visualized using heatmaps to highlight regions of interest. Areas shown in warmer colors indicate regions which the model pays more attention to. The results suggest MIRROR consistently attends to disease-relevant regions, particularly on cancerous regions. This strong alignment showcases MIRROR’s extraordinary interpretability, confirming that it not only excels in downstream tasks but also provides biologically grounded decision-making insights.
IV-E Qualitative Analysis
Figure 5: Visualization of histopathology and transcriptomics features encoded by MIRROR on the TCGA-NSCLC dataset, compared to those obtained using TANGLE. Pink dots represent samples from TCGA-LUSC, while blue dots represent samples from TCGA-LUAD. MIRROR clearly yields more distinct and representative feature distributions.
@@ IV-E Qualitative Analysis
Additionally, as depicted in Figure 5, we visualize features extracted by encoders from TANGLE and MIRROR on the TCGA-NSCLC dataset. Pink and blue dots represent TCGA-LUSC and TCGA-LUAD samples, respectively. Features from MIRROR exhibit significantly improved class separability, with minimal overlap between subtypes. In contrast, TANGLE’s feature space demonstrates considerable inter-class mixing, suggesting weaker discriminative capability. The Euclidean distances between the two modalities are also computed for each method. Notably, the distance for MIRROR is reduced by 29.02% compared to TANGLE, indicating a substantially improved alignment between modalities. These findings highlight the advantage of MIRROR in generating more representative and well-aligned feature distributions, ultimately contributing to superior model performance on downstream tasks.
V Conclusion
@@
In this paper, we introduced MIRROR, a novel multi-modal pathological SSL framework designed for joint pre-training of histopathology and transcriptomics data. Our approach is tailored to align modality-shared information while retaining modality-specific unique features. Additionally, it employs a style clustering module to reduce redundancy and preserve disease-relevant representations. Furthermore, we proposed a novel transcriptomics data preprocessing pipeline to efficiently identify disease-related genes, resulting in refined, disease-focused transcriptomics datasets. The proposed method was trained and evaluated on four distinct cohorts from the TCGA dataset, demonstrating superior performance and underscoring the efficacy of our design. Future work includes expanding evaluations to additional tasks and cohorts, as well as developing more advanced approaches for extracting information from raw transcriptomics data.

### v3

@@ IV-E Qualitative Analysis
$\lambda_{\alpha}$ $\lambda_{\beta}$ $\lambda_{\gamma}$ Subtyping Survival
0.3 0.5 0.2 0.942 ($\pm$ 0.020) 0.650 ($\pm$ 0.053)
0.3 0.2 0.5 0.950 ($\pm$ 0.021) 0.657 ($\pm$ 0.041)
0.5 0.3 0.2 0.955 ($\pm$ 0.023) 0.665 ($\pm$ 0.082)
@@
Additionally, as depicted in Figure 5, we visualize features extracted by encoders from TANGLE and MIRROR on the TCGA-NSCLC dataset. Red and blue dots represent TCGA-LUAD and TCGA-LUSC samples, respectively. Features from MIRROR exhibit significantly improved class separability, with minimal overlap between subtypes. In contrast, TANGLE’s feature space demonstrates considerable inter-class mixing, suggesting weaker discriminative capability. The Euclidean distances between the two modalities are also computed for each method. Notably, the distance for MIRROR is reduced by 29.02% compared to TANGLE, indicating a substantially improved alignment between modalities. These findings highlight the advantage of MIRROR in generating more representative and well-aligned feature distributions, ultimately contributing to superior model performance on downstream tasks.
Furthermore, as shown in Figure 6, we visualize features encoded by MIRROR and its ablated variants on the TCGA-NSCLC dataset. The variant with the alignment module alone produces compact embeddings and substantially reduces the gap between modalities, thereby achieving strong cross-modal alignment, although subtype boundaries remain indistinct. Incorporating the retention module results in more distinctive clusters and clearer separation between subtypes while preserving the improved alignment. The full model with style clustering further consolidates intra-class cohesion and enhances inter-class separation, consistent with the suppression of nuisance style redundancy and the preservation of disease-relevant information. This progression across variants highlights how the alignment, retention, and style clustering modules collectively contribute to well-aligned and highly discriminative representations.
V Discussion
V-A Gene Attribution Analysis
To gain a deeper understanding of the decision-making process of MIRROR and to provide potential biological insights into the disease, we conducted gene attribution analysis using Integrated Gradients [74]. Attribution scores were computed over the full set of input genes used during model training.
To assess biological interpretability, we examined attribution scores for genes annotated in the COSMIC database. As shown in Figure 7, many of the top-ranked genes in each cohort align well with established cancer drivers. In TCGA-BRCA, these include BRCA1/2 and ERBB2 [75, 76], in TCGA-NSCLC, EGFR [77], in TCGA-RCC, SETD2 and BAP1 [78, 79], and in TCGA-COADREAD, APC [80]. When we extended the analysis to the full training gene set, additional high-scoring features were also identified. These include PHF19 and SERPINA5 [81, 82] in TCGA-BRCA, NFATC4 [83] in TCGA-NSCLC, ARPC1B [84] in TCGA-RCC, and PHB1 [85] in TCGA-COADREAD.
These findings demonstrate the potential of MIRROR to uncover biologically meaningful patterns from multi-modal data. The use of COSMIC annotations provides a practical framework for interpretability, while future work may further substantiate these findings through external validation or functional assays.
V-B Adaptability Across Disease Types
We evaluate cross-disease adaptation performance to assess how well the learned representations extend beyond the source cohort. For both TANGLE and MIRROR, weights pretrained on TCGA-BRCA with features extracted by Phikon are frozen, and the embedding layer together with a linear classifier is trained on TCGA-NSCLC, TCGA-RCC, and TCGA-COADREAD for subtype classification and survival prediction, using the same 5-fold cross-validation splits utilized in Section IV-C. The complete results are presented in Tables IX and X. On average across the three target cohorts, MIRROR achieves a 5% improvement in subtype accuracy compared with TANGLE, and increases the C-index by 0.094 for survival prediction. These consistent gains demonstrate strong generalization beyond the source disease and confirm the superiority of MIRROR in out-of-domain transfer.
We also observe a limitation on TCGA-COADREAD, where performance reported is substantially lower than that obtained by direct pretraining and adaptation on the TCGA-COADREAD cohort. This likely reflects greater intra-cohort heterogeneity that makes transfer to TCGA-COADREAD more challenging.
Figure 5: Visualization of histopathology and transcriptomics features encoded by MIRROR on the TCGA-NSCLC dataset, compared to those obtained using TANGLE. Red dots represent samples from TCGA-LUAD, while blue dots represent samples from TCGA-LUSC. Circles denote RNA features, while triangles represent WSI features. MIRROR clearly yields more distinct and well-aligned feature distributions.
Figure 6: Visualization of histopathology and transcriptomics features encoded by MIRROR and its ablated variants on the TCGA-NSCLC dataset. Red dots represent samples from TCGA-LUAD, while blue dots represent samples from TCGA-LUSC. Circles denote RNA features, while triangles denote WSI features. The comparison clearly illustrates the contribution of each proposed module to the overall feature distributions learned by MIRROR.
Figure 7: Top 10 contributing genes identified by MIRROR across four TCGA cohorts using Integrated Gradients. Gene attribution scores were computed separately for each cohort, and the top-ranked genes were selected from those annotated in the COSMIC database. These genes highlight the most influential molecular features captured by the model in each cohort.
V-C Sensitivity Analysis
To assess the sensitivity of model predictions to small perturbations in histopathology and transcriptomic inputs, we computed two per-sample gradient-based sensitivity metrics on the TCGA-BRCA cohort.
The first sensitivity metric is the $\ell_{2}$ norm of the input gradient of the logit margin:
$s_{1}(x)\;=\;\bigl{\|}\nabla_{x}\!\bigl{(}f_{y}(x)-\max_{k\neq y}f_{k}(x)\bigr{)}\bigr{\|}_{2},$ (23)
where $f_{y}(x)$ denotes the logit of the true class label $y$ and $\max_{k\neq y}f_{k}(x)$ is the largest competing logit. This measure captures how rapidly the confidence margin between the correct class and its closest rival changes under infinitesimal input perturbations. A larger $s_{1}$ value indicates that even small feature perturbations can significantly reduce the margin and move the sample closer to a decision boundary, whereas a smaller value reflects more robust and stable predictions.
Complementary to this margin-based view, the second sensitivity metric considers the overall smoothness of the input–output mapping through the Frobenius norm of the Jacobian:
$s_{2}(x)\;=\;\bigl{\|}J(x)\bigr{\|}_{F},\qquad J(x)\;=\;\frac{\partial f(x)}{\partial x},$ (24)
where $f(x)\in\mathbb{R}^{C}$ denotes the vector of class logits for $C$ classes. The Jacobian $J(x)\in\mathbb{R}^{C\times p}$ collects the partial derivatives of each logit with respect to each input feature, with entry $J_{c,j}=\tfrac{\partial f_{c}}{\partial x_{j}}$. Here, $c\in\{1,\dots,C\}$ indexes the class dimension and $j\in\{1,\dots,p\}$ indexes the $p$ input features. The Frobenius norm provides a summary of the aggregate local sensitivity of all logits to small changes in the inputs. Larger values of $s_{2}$ imply that slight perturbations may lead to substantial variations across the output logits, whereas smaller values indicate a smoother local input–output mapping and more stable behavior [86].
Based on these two measures, MIRROR demonstrates substantially lower sensitivity than TANGLE on the TCGA-BRCA cohort, with $s_{1}=0.768\pm 0.045$ compared with $13.101\pm 3.029$ and $s_{2}=0.484\pm 0.248$ compared with $8.112\pm 4.686$. These results consistently indicate that MIRROR yields predictions that are less affected by small perturbations in histopathology or transcriptomic inputs and therefore provide greater robustness to potential preprocessing errors or feature noise.
V-D Computational Complexity
We further analyzed the computational cost of MIRROR to assess the trade-offs introduced by the proposed architecture and modules. The number of parameters and floating-point operations (FLOPs) per sample were measured and reported.
When applied to downstream tasks, where a single linear classifier is attached to the backbone, MIRROR comprises 33.1033.10M parameters and requires 90.4490.44G FLOPs per forward pass. During pretraining, the complexity increases to 48.9748.97M parameters and 139.14139.14G FLOPs, reflecting the additional overhead of these components. In comparison, TANGLE remains considerably lighter in both pretraining and downstream evaluation, with 11.4111.41M parameters and 9.699.69G FLOPs.
These results indicate that MIRROR improves robustness and multimodal integration at the expense of a larger memory footprint and greater computational demand. This trade-off highlights a limitation of the current design and points to opportunities for future work.
V-E Robustness to Domain Shift
Our experiments are conducted on four TCGA cohorts, namely TCGA-BRCA, TCGA-NSCLC, TCGA-RCC, and TCGA-COADREAD, which collectively encompass three principal sources of domain variability: institutional variation, scanner heterogeneity, and artifacts in formalin-fixed paraffin-embedded (FFPE) slides. This setup inherently evaluates the robustness of MIRROR under realistic and diverse domain shifts encountered in clinical applications.
To capture institutional variability, TCGA aggregates histopathology slides from a large network of contributing sites [87]. This multi-center composition introduces substantial heterogeneity in staining protocols, fixation methods, and tissue processing workflows. Scanner hardware also varies across sites, with both Aperio and Hamamatsu systems used in the digitization process [15], leading to differences in image resolution, color reproduction, and compression artifacts. FFPE slide preparation further introduces artifacts such as tissue folds, air bubbles, slide-edge cracks, out-of-focus regions, and pen markings. Recent quality control audits [88, 89, 90] provide quantitative evidence of the prevalence and diversity of these artifacts in TCGA cohorts.
Our 5-fold cross-validation strategy ensures that each fold includes diverse institutions, scanner types, and slide preparation modes. As a result, the reported performance reflects MIRROR’s generalization under realistic, multi-center deployment conditions.
V-F Limitations and Future Work
In this section, we reflect on the limitations of the present study and outline directions that would strengthen both performance and clinical readiness.
The current evaluation is confined to TCGA, which, although extensive and spanning multiple cancer types, does not fully capture the diversity of clinical workflows and patient populations. Expanding to external datasets will provide a more comprehensive assessment of robustness under real-world variability and yield more abundant training and evaluation samples. At the same time, pretraining is currently restricted to intra-cohort data. Extending pretraining across all TCGA cohorts and, where feasible, beyond TCGA would enable the development of a stronger and more reliable multimodal foundation model. Another limitation arises from transcriptomic inputs, which differ in length across cohorts because the preprocessing pipeline selects different gene subsets for each cohort. Establishing a harmonized gene panel across TCGA, together with the integration of pathway-level and additional omics variables using consistent preprocessing, would facilitate the training of larger and more coherent foundation models.
Beyond data-related constraints, computational efficiency also poses a challenge. The heavier backbone and modules lead to a higher parameter count and increased FLOPs compared with simpler baselines. To broaden deployability in settings with limited latency and memory budgets, future work will investigate parameter-efficient fine-tuning, knowledge distillation, and quantization in order to produce a family of models that vary in size and computational cost. In addition, practical deployments often encounter incomplete or noisy inputs, such as missing RNA profiles or variable patch-level feature quality. Addressing these issues will require modality-robust training strategies and uncertainty-aware mechanisms that can ensure reliable predictions even under imperfect input conditions.
Taken together, these directions address the principal limitations related to data breadth, modality robustness, input consistency, and computational efficiency. Advancing along these lines is expected to improve overall performance and generalization under domain shift, enhance stability and interpretability, and ultimately move MIRROR closer to serving as a clinically viable multimodal foundation model ready for deployment across diverse healthcare environments.
TABLE IX: Adaptation performance for cancer subtype classification using TCGA-BRCA pretrained weights.
Method TCGA-NSCLC TCGA-RCC TCGA-COADREAD
Acc F1 Acc F1 Acc F1
TANGLE 0.947 0.987 0.956 0.908 0.744 0.448
(± $\pm\,0.044$) (± $\pm\,0.110$) (± $\pm\,0.004$) (± $\pm\,0.148$) (± $\pm\,0.175$) (± $\pm\,0.020$)
MIRROR 0.993 0.993 0.983 0.914 0.822 0.523
(± $\pm\,0.006$) (± $\pm\,0.006$) (± $\pm\,0.012$) (± $\pm\,0.145$) (± $\pm\,0.050$) (± $\pm\,0.075$)
TABLE X: Adaptation performance for cancer survival analysis using TCGA-BRCA pretrained weights.
Method TCGA-NSCLC TCGA-RCC TCGA-COADREAD
TANGLE 0.536 (± $\pm\,0.025$) 0.744 (± $\pm\,0.130$) 0.511 (± $\pm\,0.016$)
MIRROR 0.602 (± $\pm\,0.044$) 0.778 (± $\pm\,0.055$) 0.693 (± $\pm\,0.045$)
VI Conclusion
@@
In this paper, we introduced MIRROR, a novel multi-modal pathological SSL framework designed for joint pretraining of histopathology and transcriptomics data. Our approach is tailored to align modality-shared information while retaining modality-specific unique features. Additionally, it employs a style clustering module to reduce redundancy and preserve disease-relevant representations. Furthermore, we proposed a novel transcriptomics data preprocessing pipeline to efficiently identify disease-related genes, resulting in refined, disease-focused transcriptomics datasets. The proposed method was trained and evaluated on four distinct cohorts from the TCGA dataset, demonstrating superior performance and underscoring the efficacy of our design. Future work will focus on expanding evaluations to external datasets and additional disease cohorts, extending pretraining across and beyond TCGA, and establishing harmonized molecular inputs to support larger foundation models. Efforts will also prioritize reducing computational overhead through parameter-efficient fine-tuning, knowledge distillation, and quantization, as well as developing strategies that enhance robustness to missing or noisy modalities.
Appendix A Proof of Uniqueness for the Global Minimum of the Style Loss
Proposition 1.
Let the two modality-specific posteriors be defined as:
$\mathcal{N}(\mu_{S},\sigma_{S}^{2}I),\quad\mathcal{N}(\mu_{T},\sigma_{T}^{2}I),$ (25)
with diagonal covariance in $\mathbb{R}^{D}$. The style loss:
$\begin{split}L_{\text{style}}&=\;\mathrm{KL}\bigl{(}\mathcal{N}(\mu_{S},\sigma_{S}^{2}I)\,\big{\|}\,\mathcal{N}(0,I)\bigr{)}\\ &+\,\mathrm{KL}\bigl{(}\mathcal{N}(\mu_{T},\sigma_{T}^{2}I)\,\big{\|}\,\mathcal{N}(0,I)\bigr{)},\end{split}$ (15)
has a unique global minimum:
$L_{\text{style}}^{\min}=0,$ (26)
attained if and only if:
$\mu_{S}=\mu_{T}=0,\quad\sigma_{S}=\sigma_{T}=1.$ (27)
Proof.
Consider a single latent coordinate $(\mu,\sigma)$, the KL divergence between $\mathcal{N}(\mu,\sigma^{2})$ and the standard normal distribution $\mathcal{N}(0,1)$ is given by:
$\mathrm{KL}\bigl{(}\mathcal{N}(\mu,\sigma^{2})\|\mathcal{N}(0,1)\bigr{)}=f(\mu,\sigma):=\tfrac{1}{2}\bigl{(}\mu^{2}+\sigma^{2}-1-\log\sigma^{2}\bigr{)}.$ (28)
The first–order derivatives are:
$\partial_{\mu}f=\mu,\quad\partial_{\sigma}f=\sigma-\sigma^{-1},$ (29)
which both vanish if and only if $(\mu,\sigma)=(0,1)$. Furthermore, the Hessian is given by:
$\nabla^{2}f(\mu,\sigma)=\begin{pmatrix}1&0\\[2.0pt] 0&1+\sigma^{-2}\end{pmatrix}\succ 0,$ (30)
indicating that $f$ is strictly convex. Therefore, $(0,1)$ is the unique global minimizer of $f$, with $f(0,1)=0$.
Since the KL divergence between diagonal Gaussians decomposes as a sum over coordinates, the total style loss becomes:
$L_{\text{style}}=\tfrac{1}{2}\sum_{i=1}^{D}\bigl{(}\mu_{S,i}^{2}+\sigma_{S,i}^{2}-1-\log\sigma_{S,i}^{2}+\mu_{T,i}^{2}+\sigma_{T,i}^{2}-1-\log\sigma_{T,i}^{2}\bigr{)}.$ (31)
Each summand is strictly convex and minimized uniquely at $\mu_{S,i}=\mu_{T,i}=0$ and $\sigma_{S,i}=\sigma_{T,i}=1$. Thus, the global minimum $L_{\text{style}}=0$ is achieved if and only if:
$\mu_{S}=\mu_{T}=0,\quad\sigma_{S}=\sigma_{T}=1.$ (32)
Appendix B Proof for Minimizing the Clustering Loss Increases Mutual Information
Proposition 2.
Let $\mathbf{S}_{\text{cluster}},\mathbf{T}_{\text{cluster}}\in\Delta^{K-1}$ denote the soft prototype assignments for the two modalities, and let $z_{S},z_{T}\in\mathbb{R}^{D}$ be the corresponding style representations. Define the clustering loss as:
$L_{\text{cluster}}$ $=\mathrm{KL}\!\bigl{(}\mathbf{S}_{\text{cluster}}\,\|\,\mathbf{T}_{\text{cluster}}\bigr{)}+\mathrm{KL}\!\bigl{(}\mathbf{T}_{\text{cluster}}\,\|\,\mathbf{S}_{\text{cluster}}\bigr{)}$ (19)
$=2\,\mathrm{JSD}\!\bigl{(}\mathbf{S}_{\text{cluster}},\mathbf{T}_{\text{cluster}}\bigr{)}.$
and define the function:
$g(L):=\sqrt{\frac{L}{2}}\Bigl{[}\log\!\frac{K-1}{\sqrt{L/2}}+1\Bigr{]},\quad 0<L\leq 2\log 2.$ (33)
Then, the mutual information between $z_{S}$ and $z_{T}$ satisfies:
$I(z_{S};z_{T})\;\leq\;g\bigl{(}L_{\text{cluster}}\bigr{)},$ (34)
where $g(\cdot)$ is strictly increasing. Consequently, minimizing the clustering loss $L_{\text{cluster}}$ strictly increases the mutual information shared by the two modalities.
Proof.
Since $L_{\mathrm{cluster}}\geq\mathrm{KL}(\mathbf{S}_{\text{cluster}}\|\mathbf{T}_{\text{cluster}})$, Pinsker’s inequality [91] gives:
$\|\mathbf{S}_{\text{cluster}}-\mathbf{T}_{\text{cluster}}\|_{1}\leq\sqrt{2L_{\mathrm{cluster}}}.$ (35)
To relate this bound to a mismatch probability, consider the two categorical distributions $\mathbf{S}_{\text{cluster}}=(s_{1},\dots,s_{K})$ and $\mathbf{T}_{\text{cluster}}=(t_{1},\dots,t_{K})$. For each index $k$, define:
$\begin{split}r_{k}&:=\min(s_{k},t_{k}),\\ a_{k}&:=\max(s_{k}-t_{k},0),\\ b_{k}&:=\max(t_{k}-s_{k},0),\end{split}$ (36)
so that $r_{k}$ is the shared mass, and $a_{k}$ and $b_{k}$ are the excess contributions from Slide and Transcriptomics modalities, respectively. Let $r:=\sum_{k}r_{k}$, then $\sum_{k}a_{k}=\sum_{k}b_{k}=1-r$ and $\lVert\mathbf{S}_{\text{cluster}}-\mathbf{T}_{\text{cluster}}\rVert_{1}=2(1-r)$.
Now construct a coupling $C^{\ast}=(c_{kl}^{\ast})$ as follows: assign the shared mass to the diagonal, $c_{kk}^{\ast}=r_{k}$. Since the total surplus is equal on both sides, define a bijection $\pi:\{k:a_{k}>0\}\to\{k:b_{k}>0\}$, and let:
$c^{*}_{k,\pi(k)}=\tfrac{1}{2}a_{k},\quad c^{*}_{\pi(k),k}=\tfrac{1}{2}b_{\pi(k)}.$ (37)
Set all remaining entries to zero. By construction, $C^{\ast}$ is a valid joint distribution with marginals $\mathbf{S}_{\mathrm{cluster}}$ and $\mathbf{T}_{\mathrm{cluster}}$.
Under this coupling, disagreement occurs only in the surplus mass. Hence, the probability of mismatch is:
$\begin{split}p_{e}&:=\Pr_{C^{\ast}}[K_{S}\neq K_{T}]\\ &=1-r\\ &=\tfrac{1}{2}\lVert\mathbf{S}_{\text{cluster}}-\mathbf{T}_{\text{cluster}}\rVert_{1}\\ &=d_{\mathrm{TV}}\bigl{(}\mathbf{S}_{\text{cluster}},\mathbf{T}_{\text{cluster}}\bigr{)}.\end{split}$ (38)
Combining Eq. (38) with Eq. (35), we have:
$p_{e}\;\leq\;\sqrt{\tfrac{L_{\mathrm{cluster}}}{2}}.$ (39)
Applying Fano’s inequality [92], we have:
$\begin{split}H(K_{S}\!\mid K_{T})&\leq h_{2}(p_{e})+p_{e}\log(K-1)\\ &\leq p_{e}\Bigl{[}\log\tfrac{K-1}{p_{e}}+1\Bigr{]},\end{split}$ (40)
where $h_{2}(\cdot)$ is the binary entropy function. Substituting the bound Eq. (39), we obtain:
$H(K_{S}\!\mid K_{T})\leq g\!\bigl{(}L_{\mathrm{cluster}}\bigr{)}.$ (41)
Since $I(K_{S};K_{T})=H(K_{S})-H(K_{S}\!\mid K_{T})$ and $H(K_{S})\geq 0$, we have:
$I(K_{S};K_{T})\;\leq\;g\!\bigl{(}L_{\mathrm{cluster}}\bigr{)}.$ (42)
Note that $L_{\mathrm{cluster}}=2\,\mathrm{JSD}(\mathbf{S}_{\mathrm{cluster}},\mathbf{T}_{\mathrm{cluster}})$, and since $\mathrm{JSD}(P,Q)\leq\log 2$ for any pair of distributions $P,Q$, it follows that:
$0<L_{\mathrm{cluster}}\leq 2\log 2.$ (43)
This justifies the domain restriction in the definition of $g(L)$. Now observe that:
$g^{\prime}(L)=\frac{1}{2\sqrt{2L}}\left[\log(K-1)+\tfrac{1}{2}\log 2-\tfrac{1}{2}\log L\right]>0$ (44)
for $0<L\leq 2\log 2$, so $g(\cdot)$ is strictly increasing.
Because the prototype assignments $K_{S}$ and $K_{T}$ are deterministic functions of the latents $z_{S}$ and $z_{T}$, the data processing inequality [93] implies:
$I(K_{S};K_{T})\leq I(z_{S};z_{T}).$ (45)
Taking negatives and combining with the previous bound yields:
$I(z_{S};z_{T})\;\leq\;-\;I(K_{S};K_{T})\;\leq\;g\!\bigl{(}L_{\mathrm{cluster}}\bigr{)}.$ (46)
Since $g(\cdot)$ is strictly increasing, reducing $L_{\text{cluster}}$ strictly reduces the upper bound, thereby increasing the mutual information $I(z_{S};z_{T})$. ∎